OpenStudy (anonymous):

Find the equation of the line passing through the point (x,y)=(4,-2), which is parallel to the line y= -x-7

OpenStudy (yttrium):

Do know how to get the slope of the given line? If you know, just get it and substitute the values of x and y to this from: Y-Y1 = m (x-x1)

OpenStudy (anonymous):

y=-x+k then substitute point u'll get

OpenStudy (anonymous):

got the answer??

OpenStudy (anonymous):

You here?

OpenStudy (anonymous):

respond respond!!!

OpenStudy (anonymous):

Thanks for the tips. I managed to figure it out.